OpenStudy (anonymous):

what is the molarity of a solution in which 3 moles of AgNO3 is dissolve in 1.5 liter of water?

OpenStudy (aaronq):

use the formula \(M=\dfrac{n_{solute}}{L_{solution}}\)

OpenStudy (anonymous):

can you show me ?

OpenStudy (anonymous):

step by step?

OpenStudy (anonymous):

M=n/v M=molarity n=number of moles v=volume in liters so... M=3moles AgNO3/1.5 liters M=2

OpenStudy (anonymous):

ok and than?

OpenStudy (anonymous):

M stands for molarity. Therefore, your answer is 2.